OpenStudy (anonymous):

Evaluate the limit as x approaches 1 of the quotient of the square root of the quantity x squared plus 8 minus 3 and the quantity x minus 1. You must show your work to receive credit.

OpenStudy (anonymous):

The two-sided limit does not exist. limit as x approaches 1 of [(square root x^2+8) -3]/[x-1] limit as x approaches 1 of [(square root 1^2+8)-3]/[1-1] limit as x approaches 1 of [(square root 9)-3]/[0] limit as x approaches 1 of [(3)-(3)]/[0] limit as x approaches 1 of [0]/[0] DNE, undefined.

OpenStudy (anonymous):

@Luigi0210 please help!! :(

OpenStudy (anonymous):

@Luigi0210 is there another way to evaluate this limit or is this the only way? could we find a way to not make the bottom 0?

OpenStudy (luigi0210):

Do you know L'hopital's ?

OpenStudy (anonymous):

@ikram002p @Luigi0210 no can u please write out the steps so i can follow them :) this is an essay question so i neeed to understand every step of the problem.

OpenStudy (luigi0210):

Calc I right?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

are you there ? :)

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

If you don't know it then I assume they don't want you to know about it yet. So yea, when it approaches 1, it is DNE

OpenStudy (anonymous):

no i think im supposed 2 know it

OpenStudy (anonymous):

i just need 2 brush up on it :)

OpenStudy (anonymous):

i know were gonna have 2 factor? i think

OpenStudy (luigi0210):

Pretty much this: http://prntscr.com/65l363 If you get 0/0, or inf/inf, then you are allowed to take the derivatives of both and try the limit again

OpenStudy (anonymous):

ok thanks can u show the steps on here plz? :)

OpenStudy (luigi0210):

Okay so we have \(\Large \lim_{x \rightarrow 1} ~\frac{\sqrt{x^2+8}}{x-1} \) which turned out to be \(\frac{0}{0} \) . Now, according to L'hopital's, we can do this: \(\Large \lim_{x \rightarrow 1} ~\frac{dy/dx (\sqrt{x^2+8})}{dy/dx (x-1)} \)

OpenStudy (luigi0210):

Do you know how to take the derivatives?

OpenStudy (luigi0210):

Or do you need help on that?

OpenStudy (anonymous):

yes i need help on it :) btw thanks for the 1st step :)

OpenStudy (luigi0210):

So for \(\large \sqrt{x^2+8} \) can be rewritten as \(\large (x^2+8)^{1/2} \) . For this one we have to use chain rule: First take the derivative of the outside, which is bringing down the 1/2 and subtracting 1. \(\Large \color{red}{\frac{1}{2}} * (x^2+8)^{\color{red}{\frac{1}{2}-1}} \) We get \(\Large \frac{1}{2}(x^2+8)^{\color{green}{-1/2}} \) Now for the chain rule, take the derivative of the inside: \(\Large \frac{1}{2} (\color{blue}{x^2+8})^{-1/2} *\color{blue}{2x+0} \) Since dy/dx of x^2 is 2x, because we bring down the 2 and subtract 1. And the dy/dx of any constant is 0. Make sense so far?

OpenStudy (anonymous):

yes :)!!! perfect explanation

OpenStudy (luigi0210):

Okay good. Now, if you remember the rules of exponents, whenever it is negative, we are allowed to flip it to make it positive, so it becomes: \(\Large \frac{2x}{2\sqrt{x^2+8}} \) The 2's can cancel: \(\Large \frac{x}{\sqrt{x^2+8}} \) And we have that^

OpenStudy (luigi0210):

Now, think you can take the dy/dx of x-1?

OpenStudy (anonymous):

wouldn't it just be 0?

OpenStudy (anonymous):

actually the answer is 1 because of the sum rule...right?

OpenStudy (luigi0210):

Nopee, remember the x still have a power of 1. So it would actually be: \(\large \color{red}{1} x^{\color{red}{1-1}} =1*(x^0)\) and x^0 is 1. so: \(\Large 1*1=1 \) The -1 however is a constant, so that is 0

OpenStudy (freckles):

Without LH: \[\lim_{x \rightarrow 1}\frac{\sqrt{x^2+8}-3}{x-1} \\ \text{ you can also choose to do this without derivatives } \\ \text{ you can decide to rationalize the numerator by } \\ \text{ multiplying top and bottom by top's conjugate } \\ \lim_{x \rightarrow 1}\frac{\sqrt{x^2+8}-3}{x-1} \cdot \frac{\sqrt{x^2+8}+3}{\sqrt{x^2+8}+3} \\ \lim_{x \rightarrow 1}\frac{(x^2+8)-9}{(x-1) (\sqrt{x^2+8}+3)} \] \[\text{ You will need to know how to factor a difference of squares } \\ x^2-a^2=(x-a)(x+a)\]

OpenStudy (luigi0210):

Yea, the conjugate works too.. depends which way you wanna go ^

OpenStudy (anonymous):

ohh freckles i like that idea!!! it seems easier, luigi do we have a long way 2 go?

OpenStudy (luigi0210):

Not really, now all we have is: \(\Large \lim_{x \rightarrow 1} \frac{x}{\sqrt {x^2+8}} -3 \) :P

OpenStudy (freckles):

hey what's that -3 there for?

OpenStudy (luigi0210):

Oh, whoops, yea, sorry about that xD

OpenStudy (luigi0210):

Ignore that -3. it turns to 0

OpenStudy (anonymous):

ok freckles can u please show the next step after limit as x approaches 1 of (x^2+8)-9/(x-1)[(square root x^2+8) +3]

OpenStudy (anonymous):

okay then luigi do i just plug in the 1 now?

OpenStudy (luigi0210):

Yup, it should give you an answer.. if not, well then L'Hop again if it's 0/0 xD

OpenStudy (anonymous):

i think it gives: 1/3

OpenStudy (anonymous):

is it correct?:)

OpenStudy (luigi0210):

It most definitely is :P

OpenStudy (anonymous):

THANKU!!!!!!!!!!!!!!!!!!!!! LIFE SAVERRRRRRRR

OpenStudy (anonymous):

i have 6 more questions, are you available to help? :)

OpenStudy (luigi0210):

You're welcome xD And sure, I can try :P

OpenStudy (freckles):

I know you got the answer...But you asked what happens after that Well the hint was to fact the difference of squares. The difference of squares was on top if you simplified the top. That is the top should be x^2-1 which equals (x-1)(x+1) Now you will be able to cancel a common factor from top and bottom That is you should have: \[ \lim_{x \rightarrow 1}\frac{\sqrt{x^2+8}-3}{x-1} \\ \\ \text{ you can decide to rationalize the numerator by } \\ \text{ multiplying top and bottom by top's conjugate } \\ \lim_{x \rightarrow 1}\frac{\sqrt{x^2+8}-3}{x-1} \cdot \frac{\sqrt{x^2+8}+3}{\sqrt{x^2+8}+3} \\ \lim_{x \rightarrow 1}\frac{(x^2+8)-9}{(x-1) (\sqrt{x^2+8}+3)} \\ \lim_{x \rightarrow 1} \frac{x^2-1}{(x-1)(\sqrt{x^2+8}+3)} \\ \lim_{ x \rightarrow 1}\frac{(x-1)(x+1)}{(x-1)(\sqrt{x^2+8}+3)} \\ \lim_{x \rightarrow 1} \frac{x+1}{\sqrt{x^2+8}+3} \] Now you can replace x with 1.